A spinner has 10 equally sized sections, 7 of which are gray and 3 of which are blue. The spinner is spun twice. What is the probability that the first spin lands on gray and the second spin lands on blue ?

Write your answer as a fraction in simplest form.
Answer #1

Here we are given 7 are gray and 3 are blue so

P(first on gray then blue)=(7*3/10*10)=21/100